# Zero-Shot Learning of Causal Models

### v2

@@ Appendix B Experiments on Real World Benchmark
Appendix C Results without True Causal Graph
@@
Our results in the main paper (Table 1Table 2Table 3) require the knowledge of true graph (${\bm{\mathcal{G}}}$) as part of the input context to Cond-FiP. In this section we conduct where we don’t provide the true graph in the input context, rather we infer the graph $\hat{\bm{\mathcal{G}}}$ using an amortized causal discovery approach (AVICI (Lorch et al., 2022)) from the observational data $D_{\bm{X}}$. We chose AVICI for this task since it can enable to infer the graph in a zero-shot manner, hence allowing the combined pipeline of AVICI + Cond-FiP to be a zero-shot technique for learning SCMs. More precisely, AVICI zero-shot infers the graph ${\bm{\mathcal{G}}}$ from input context $D_{\bm{X}}$, and we pass $(\hat{\bm{\mathcal{G}}},D_{\bm{X}})$ as the input context for Cond-FiP. Therefore, for any $\bm{z}\in\mathbb{R}^{d}$, Cond-FiP ( ${\mathcal{T}}(\bm{z},D_{\bm{X}},\hat{\bm{\mathcal{G}}})$) aims to replicate the functional mechanism ${\bm{F}}({\bm{z}})$ of the underlying SCM.
@@
The results for benchmarking Cond-FiP with inferred graphs using AVICI for the task of noise prediction, sample generation, and interventional generation are provided in Table 5Table 6, and Table 7 respectively. For a fair comparison, the baselines FiP, DECI, and DoWhy also use the inferred graph ($\hat{\bm{\mathcal{G}}}$) by AVICI instead of the true graph (${\bm{\mathcal{G}}}$). We find that Cond-FiP remains competitive to baselines even for the scenario of unknown true causal graph. Hence, our amortized training procedure can be extended easily for zero-shot inference of both graphs and causal mechanisms of the SCM.
@@
Table 5: Results for Noise Prediction without True Graph. We compare Cond-FiP against the baselines for the task of predicting noise variable from input observations. Unlike experiments in the main paper, we do not use the true graph, and infer the graph first in a zero-shot manner using AVICI (Lorch et al., 2022). Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP.
@@
Table 6: Results for Sample Generation without True Graph. We compare Cond-FiP against the baselines for the task of generating samples from the input noise variable. Unlike experiments in the main paper, we do not use the true graph, and infer the graph first in a zero-shot manner using AVICI (Lorch et al., 2022). Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP.
@@
Table 7: Results for Interventional Generation without True Graph. We compare Cond-FiP against the baselines for the task of interventional data from the input noise variable. Unlike experiments in the main paper, we do not use the true graph, and infer the graph first in a zero-shot manner using AVICI (Lorch et al., 2022). Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP.
Appendix D Evaluating Generalization of Cond-Fip in Scarce Data Regime
D.1 Experiments with $n_{\textnormal{test}}=100$
In this section we benchmark Cond-FiP against the baselines for the scenario when test datasets in the input context have smaller sample size ($n_{\textnormal{test}}=100$) as compared to the train datasets ($n_{\textnormal{test}}=400$).
We report the results for the task of noise prediction, sample generation, and interventional generation in Table 8Table 9, and Table 10 respectively. We find that Cond-FiP exhibits superior generalization as compared to baselines. For example, in the case of RFF IN, Cond-FiP is even better than FiP for all the tasks! This can be attributed to the advantage of zero-shot inference; as the sample size in test dataset decreases, the generalization of baselines would be affected a lot since they require training from scratch on these datasets. However, zero-shot inference methods would be impacted less as they do not have to trained from scratch, and the inductive bias learned by them can help them generalize even with smaller input context.
Method Total Nodes LIN IN RFF IN LIN OUT RFF OUT
DoWhy 10 $0.06$ $(0.01)$ $0.22$ $(0.03)$ $0.09$ $(0.01)$ $0.16$ $(0.03)$
DECI 10 $0.15$ $(0.01)$ $0.3$ $(0.02)$ $0.22$ $(0.01)$ $0.3$ $(0.03)$
FiP 10 $0.07$ $(0.01)$ $0.18$ $(0.01)$ $0.12$ $(0.01)$ $0.11$ $(0.01)$
Cond-FiP 10 $0.07$ $(0.01)$ $0.14$ $(0.01)$ $0.09$ $(0.01)$ $0.14$ $(0.01)$
DoWhy 20 $0.06$ $(0.01)$ $0.27$ $(0.05)$ $0.07$ $(0.01)$ $0.37$ $(0.01)$
DECI 20 $0.15$ $(0.02)$ $0.33$ $(0.02)$ $0.17$ $(0.02)$ $0.35$ $(0.03)$
FiP 20 $0.09$ $(0.01)$ $0.21$ $(0.03)$ $0.1$ $(0.01)$ $0.27$ $(0.03)$
Cond-FiP 20 $0.08$ $(0.01)$ $0.12$ $(0.01)$ $0.1$ $(0.01)$ $0.15$ $(0.01)$
DoWhy 50 $0.06$ $(0.01)$ $0.29$ $(0.04)$ $0.05$ $(0.01)$ $0.47$ $(0.04)$
DECI 50 $0.14$ $(0.01)$ $0.33$ $(0.02)$ $0.14$ $(0.02)$ $0.4$ $(0.03)$
FiP 50 $0.08$ $(0.01)$ $0.23$ $(0.03)$ $0.08$ $(0.01)$ $0.37$ $(0.04)$
Cond-FiP 50 $0.08$ $(0.0)$ $0.12$ $(0.01)$ $0.08$ $(0.01)$ $0.15$ $(0.01)$
DoWhy 100 $0.06$ $(0.01)$ $0.31$ $(0.04)$ $0.06$ $(0.01)$ $0.5$ $(0.03)$
DECI 100 $0.13$ $(0.01)$ $0.36$ $(0.03)$ $0.12$ $(0.02)$ $0.44$ $(0.02)$
FiP 100 $0.08$ $(0.01)$ $0.25$ $(0.04)$ $0.1$ $(0.01)$ $0.39$ $(0.03)$
Cond-FiP 100 $0.07$ $(0.0)$ $0.13$ $(0.01)$ $0.08$ $(0.01)$ $0.17$ $(0.01)$
Table 8: Results for Noise Prediction with Smaller Sample Size ($n_{\textnormal{test}}=100$). We compare Cond-FiP against the baselines for the task of predicting noise variable from input observations. Each test dataset contains $100$ samples, as opposed to $400$ samples in the main paper. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP.
Method Total Nodes LIN IN RFF IN LIN OUT RFF OUT
DoWhy 10 $0.1$ $(0.01)$ $0.3$ $(0.06)$ $0.12$ $(0.02)$ $0.19$ $(0.03)$
DECI 10 $0.23$ $(0.01)$ $0.45$ $(0.04)$ $0.31$ $(0.02)$ $0.38$ $(0.04)$
FiP 10 $0.13$ $(0.01)$ $0.29$ $(0.04)$ $0.18$ $(0.02)$ $0.15$ $(0.03)$
Cond-FiP 10 $0.09$ $(0.01)$ $0.2$ $(0.03)$ $0.09$ $(0.02)$ $0.14$ $(0.02)$
DoWhy 20 $0.11$ $(0.01)$ $0.47$ $(0.15)$ $0.11$ $(0.02)$ $0.5$ $(0.03)$
DECI 20 $0.26$ $(0.02)$ $0.53$ $(0.05)$ $0.26$ $(0.03)$ $0.57$ $(0.04)$
FiP 20 $0.17$ $(0.02)$ $0.34$ $(0.06)$ $0.17$ $(0.02)$ $0.39$ $(0.03)$
Cond-FiP 20 $0.08$ $(0.0)$ $0.31$ $(0.06)$ $0.13$ $(0.01)$ $0.37$ $(0.02)$
DoWhy 50 $0.11$ $(0.01)$ $0.42$ $(0.08)$ $0.09$ $(0.01)$ $0.66$ $(0.06)$
DECI 50 $0.23$ $(0.02)$ $0.59$ $(0.08)$ $0.27$ $(0.04)$ $0.73$ $(0.06)$
FiP 50 $0.13$ $(0.01)$ $0.38$ $(0.07)$ $0.14$ $(0.01)$ $0.58$ $(0.06)$
Cond-FiP 50 $0.1$ $(0.01)$ $0.32$ $(0.05)$ $0.12$ $(0.01)$ $0.54$ $(0.05)$
DoWhy 100 $0.11$ $(0.01)$ $0.44$ $(0.08)$ $0.11$ $(0.01)$ $0.74$ $(0.05)$
DECI 100 $0.25$ $(0.02)$ $0.62$ $(0.08)$ $0.25$ $(0.01)$ $0.78$ $(0.07)$
FiP 100 $0.15$ $(0.01)$ $0.4$ $(0.07)$ $0.19$ $(0.02)$ $0.67$ $(0.07)$
Cond-FiP 100 $0.11$ $(0.01)$ $0.35$ $(0.07)$ $0.14$ $(0.02)$ $0.63$ $(0.07)$
Table 9: Results for Sample Generation with Smaller Sample Size ($n_{\textnormal{test}}=100$). We compare Cond-FiP against the baselines for the task of generating samples from the input noise variable. Each test dataset contains $100$ samples, as opposed to $400$ samples in the main paper. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP.
Method Total Nodes LIN IN RFF IN LIN OUT RFF OUT
DoWhy 10 $0.09$ $(0.01)$ $0.34$ $(0.08)$ $0.11$ $(0.01)$ $0.2$ $(0.04)$
DECI 10 $0.24$ $(0.02)$ $0.43$ $(0.04)$ $0.26$ $(0.03)$ $0.35$ $(0.04)$
FiP 10 $0.13$ $(0.01)$ $0.29$ $(0.04)$ $0.14$ $(0.02)$ $0.14$ $(0.03)$
Cond-FiP 10 $0.09$ $(0.02)$ $0.21$ $(0.03)$ $0.09$ $(0.01)$ $0.12$ $(0.02)$
DoWhy 20 $0.1$ $(0.01)$ $0.37$ $(0.08)$ $0.11$ $(0.02)$ $0.49$ $(0.04)$
DECI 20 $0.25$ $(0.03)$ $0.5$ $(0.05)$ $0.28$ $(0.03)$ $0.54$ $(0.04)$
FiP 20 $0.16$ $(0.01)$ $0.33$ $(0.06)$ $0.2$ $(0.03)$ $0.38$ $(0.03)$
Cond-FiP 20 $0.1$ $(0.01)$ $0.27$ $(0.05)$ $0.15$ $(0.02)$ $0.29$ $(0.03)$
DoWhy 50 $0.12$ $(0.02)$ $0.49$ $(0.14)$ $0.09$ $(0.01)$ $0.64$ $(0.07)$
DECI 50 $0.26$ $(0.03)$ $0.56$ $(0.07)$ $0.26$ $(0.03)$ $0.72$ $(0.06)$
FiP 50 $0.16$ $(0.02)$ $0.36$ $(0.06)$ $0.15$ $(0.01)$ $0.57$ $(0.06)$
Cond-FiP 50 $0.13$ $(0.02)$ $0.29$ $(0.04)$ $0.12$ $(0.01)$ $0.49$ $(0.07)$
DoWhy 100 $0.11$ $(0.01)$ $0.46$ $(0.07)$ $0.11$ $(0.01)$ $1.16$ $(0.38)$
DECI 100 $0.24$ $(0.02)$ $0.62$ $(0.08)$ $0.26$ $(0.01)$ $0.78$ $(0.07)$
FiP 100 $0.16$ $(0.02)$ $0.39$ $(0.07)$ $0.2$ $(0.02)$ $0.66$ $(0.07)$
Cond-FiP 100 $0.12$ $(0.02)$ $0.32$ $(0.07)$ $0.13$ $(0.01)$ $0.58$ $(0.07)$
Table 10: Results for Interventional Generation with Smaller Sample Size ($n_{\textnormal{test}}=100$). We compare Cond-FiP against the baselines for the task of generating interventional data from the input noise variable. Each test dataset contains $100$ samples, as opposed to $400$ samples in the main paper. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP.
D.2 Experiments with $n_{\textnormal{test}}=50$
We conduct more experiments for the smaller sample size scenarios, where decrease the sample size even further to $n_{\textnormal{test}}=50$ samples. We report the results for the task of noise prediction, sample generation, and interventional generation in Table 11Table 12, and Table 13 respectively. We find that baselines perform much worse than Cond-FiP for the all different SCM distributions, highlighting the efficacy of Cond-FiP for inferring functional relationships when the input context has smaller sample size. Note that there were issues with training DoWhy for such a small dataset, hence we do not consider them for this scenario.
Method Total Nodes LIN IN RFF IN LIN OUT RFF OUT
DECI 10 $0.19$ $(0.02)$ $0.41$ $(0.03)$ $0.2$ $(0.02)$ $0.42$ $(0.04)$
FiP 10 $0.13$ $(0.03)$ $0.27$ $(0.03)$ $0.15$ $(0.02)$ $0.21$ $(0.03)$
Cond-FiP 10 $0.09$ $(0.01)$ $0.17$ $(0.01)$ $0.11$ $(0.01)$ $0.16$ $(0.01)$
DECI 20 $0.2$ $(0.01)$ $0.42$ $(0.03)$ $0.25$ $(0.04)$ $0.45$ $(0.05)$
FiP 20 $0.12$ $(0.01)$ $0.33$ $(0.04)$ $0.15$ $(0.02)$ $0.35$ $(0.04)$
Cond-FiP 20 $0.1$ $(0.01)$ $0.16$ $(0.01)$ $0.11$ $(0.01)$ $0.17$ $(0.01)$
DECI 50 $0.2$ $(0.02)$ $0.43$ $(0.02)$ $0.2$ $(0.03)$ $0.5$ $(0.05)$
FiP 50 $0.13$ $(0.01)$ $0.32$ $(0.03)$ $0.13$ $(0.01)$ $0.49$ $(0.05)$
Cond-FiP 50 $0.1$ $(0.01)$ $0.16$ $(0.0)$ $0.1$ $(0.01)$ $0.17$ $(0.01)$
DECI 100 $0.19$ $(0.02)$ $0.43$ $(0.03)$ $0.21$ $(0.01)$ $0.53$ $(0.02)$
FiP 100 $0.11$ $(0.01)$ $0.32$ $(0.04)$ $0.13$ $(0.01)$ $0.48$ $(0.02)$
Cond-FiP 100 $0.09$ $(0.01)$ $0.16$ $(0.01)$ $0.09$ $(0.01)$ $0.18$ $(0.01)$
Table 11: Results for Noise Prediction with Smaller Sample Size ($n_{\textnormal{test}}=50$). We compare Cond-FiP against the baselines for the task of predicting noise variable from input observations. Each test dataset contains $50$ samples, as opposed to $400$ samples in the main paper. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP.
Method Total Nodes LIN IN RFF IN LIN OUT RFF OUT
DECI 10 $0.31$ $(0.02)$ $0.58$ $(0.05)$ $0.27$ $(0.04)$ $0.49$ $(0.07)$
FiP 10 $0.2$ $(0.03)$ $0.4$ $(0.05)$ $0.21$ $(0.03)$ $0.25$ $(0.04)$
Cond-FiP 10 $0.12$ $(0.02)$ $0.28$ $(0.03)$ $0.12$ $(0.01)$ $0.18$ $(0.03)$
DECI 20 $0.34$ $(0.02)$ $0.66$ $(0.08)$ $0.39$ $(0.07)$ $0.68$ $(0.05)$
FiP 20 $0.2$ $(0.01)$ $0.51$ $(0.08)$ $0.25$ $(0.04)$ $0.51$ $(0.02)$
Cond-FiP 20 $0.13$ $(0.01)$ $0.4$ $(0.06)$ $0.19$ $(0.02)$ $0.43$ $(0.02)$
DECI 50 $0.32$ $(0.02)$ $0.66$ $(0.06)$ $0.36$ $(0.02)$ $0.8$ $(0.06)$
FiP 50 $0.2$ $(0.01)$ $0.48$ $(0.07)$ $0.22$ $(0.02)$ $0.69$ $(0.06)$
Cond-FiP 50 $0.15$ $(0.02)$ $0.4$ $(0.05)$ $0.16$ $(0.01)$ $0.59$ $(0.06)$
DECI 100 $0.36$ $(0.04)$ $0.68$ $(0.08)$ $0.39$ $(0.03)$ $0.84$ $(0.06)$
FiP 100 $0.2$ $(0.02)$ $0.49$ $(0.09)$ $0.28$ $(0.03)$ $0.73$ $(0.07)$
Cond-FiP 100 $0.16$ $(0.01)$ $0.42$ $(0.07)$ $0.22$ $(0.01)$ $0.68$ $(0.06)$
Table 12: Results for Sample Generation with Smaller Sample Size ($n_{\textnormal{test}}=50$). We compare Cond-FiP against the baselines for the task of generating samples from the input noise variable. Each test dataset contains $50$ samples, as opposed to $400$ samples in the main paper. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP.
Method Total Nodes LIN IN RFF IN LIN OUT RFF OUT
DECI 10 $0.3$ $(0.03)$ $0.53$ $(0.05)$ $0.26$ $(0.04)$ $0.42$ $(0.05)$
FiP 10 $0.21$ $(0.04)$ $0.35$ $(0.04)$ $0.2$ $(0.03)$ $0.22$ $(0.03)$
Cond-FiP 10 $0.12$ $(0.01)$ $0.19$ $(0.03)$ $0.07$ $(0.01)$ $0.14$ $(0.02)$
DECI 20 $0.33$ $(0.02)$ $0.6$ $(0.06)$ $0.43$ $(0.07)$ $0.63$ $(0.04)$
FiP 20 $0.21$ $(0.02)$ $0.46$ $(0.07)$ $0.29$ $(0.04)$ $0.49$ $(0.02)$
Cond-FiP 20 $0.11$ $(0.01)$ $0.29$ $(0.06)$ $0.15$ $(0.02)$ $0.32$ $(0.03)$
DECI 50 $0.34$ $(0.02)$ $0.66$ $(0.07)$ $0.34$ $(0.02)$ $0.78$ $(0.06)$
FiP 50 $0.21$ $(0.02)$ $0.46$ $(0.07)$ $0.23$ $(0.02)$ $0.68$ $(0.06)$
Cond-FiP 50 $0.13$ $(0.02)$ $0.31$ $(0.05)$ $0.12$ $(0.02)$ $0.51$ $(0.07)$
DECI 100 $0.37$ $(0.04)$ $0.67$ $(0.08)$ $0.4$ $(0.04)$ $0.84$ $(0.06)$
FiP 100 $0.21$ $(0.02)$ $0.49$ $(0.08)$ $0.28$ $(0.03)$ $0.73$ $(0.07)$
Cond-FiP 100 $0.12$ $(0.01)$ $0.33$ $(0.07)$ $0.14$ $(0.01)$ $0.58$ $(0.07)$
Table 13: Results for Interventional Generation with Smaller Sample Size ($n_{\textnormal{test}}=50$). We compare Cond-FiP against the baselines for the task of generating interventional data from the input noise variable. Each test dataset contains $50$ samples, as opposed to $400$ samples in the main paper. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP.
Appendix E Ablation Study
E.1 Ablation Study of Encoder
Similar to the ablation study of decoder Appendix E, we conduct an ablation study where we train two variants of the encoder in Cond-FiP described as follows:
  • Cond-FiP (LIN): We sample SCMs with linear functional relationships during training of the encoder.
  • Cond-FiP (RFF): We sample SCMs with rff functional relationships during training of the encoder.
Note that for the training the subsequent decoder, we sample SCMs with both linear and rff functional relationships as in the main results ( Table 1Table 2, and Table 3). Note that in the main results, the encoder was trained by sampling SCMs with both linear and rff functional relationships. Hence, this ablation helps us to understand whether the strategy of training encoder on mixed functional relationships can bring more generalization to the amortization process, or if we should have trained encoders specialized for linear and non-linear functional relationships.
We present our results of the ablation study for the task of noise prediction, sample generation, and interventional generation in Table 14Table 15Table 16 respectively. Our findings indicate that Cond-FiP is robust to the choice of encoder training strategy! Even though the encoder for Cond-FiP (LIN) was only trained on data from linear SCMs, its generalization performance is similar to Cond-FiP where the encoder was trained on data from both linear and non-linear SCMs.

### v3

@@ E.2 Experiments with $n_{\mathcal{D_{\mathrm{test}}}}=50$
Appendix F Experiments without True Causal Graph on AVICI Benchmark
@@
Results in Appendix C (Table 1Table 2Table 3) require the knowledge of true graph (${\bm{\mathcal{G}}}$) as part of the input context to Cond-FiP. In this section we conduct where we don’t provide the true graph in the input context, rather we infer the graph $\hat{\bm{\mathcal{G}}}$ using an amortized causal discovery approach (AVICI [Lorch et al., 2022]) from the observational data $D_{\bm{X}}$. We chose AVICI for this task since it can enable to amortized inference of causal graphs, hence allowing the combined pipeline of AVICI + Cond-FiP can perform amortized inference of SCMs. More precisely, AVICI infers the graph from a novel instance ${\bm{\mathcal{G}}}$ from input context $D_{\bm{X}}$ without updating any parameters, and we pass $(\hat{\bm{\mathcal{G}}},D_{\bm{X}})$ as the input context for Cond-FiP. Therefore, for any $\bm{z}\in\mathbb{R}^{d}$, Cond-FiP ( ${\mathcal{T}}(\bm{z},D_{\bm{X}},\hat{\bm{\mathcal{G}}})$) aims to replicate the functional mechanism ${\bm{F}}({\bm{z}})$ of the underlying SCM.
@@
The results for benchmarking Cond-FiP with inferred graphs using AVICI for the task of noise prediction, sample generation, and interventional generation are provided in Table 16Table 17, and Table 18 respectively. For a fair comparison, the baselines FiP, DECI, and DoWhy also use the inferred graph ($\hat{\bm{\mathcal{G}}}$) by AVICI instead of the true graph (${\bm{\mathcal{G}}}$). We find that Cond-FiP remains competitive to baselines even for the scenario of unknown true causal graph. Hence, our training procedure can be extended for amortized inference of both causal graphs and causal mechanisms of the SCM.
@@
Table 16: Results for Noise Prediction without True Graph. We compare Cond-FiP against the baselines for the task of predicting noise variable from input observations. Unlike experiments in Table 1, the true graph ${\bm{\mathcal{G}}}$ is not present in input context, rather its inferred via AVICI [Lorch et al., 2022]. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results indicate Cond-FiP can generalize to novel instances even in the absence of true graph.
@@
Table 17: Results for Sample Generation without True Graph. We compare Cond-FiP against the baselines for the task of generating samples from the input noise variable. Unlike experiments in Table 2, the true graph ${\bm{\mathcal{G}}}$ is not present in input context, rather its inferred via AVICI [Lorch et al., 2022].. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results indicate Cond-FiP can generalize to novel instances even in the absence of true graph.
@@
Table 18: Results for Interventional Generation without True Graph. We compare Cond-FiP against the baselines for the task of interventional data from the input noise variable. Unlike experiments in Table 3, the true graph ${\bm{\mathcal{G}}}$ is not present in input context, rather its inferred via AVICI [Lorch et al., 2022]. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows deonte the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results indicate Cond-FiP can generalize to novel instances even in the absence of true graph.
Appendix G Ablation Study on AVICI benchmark
G.1 Ablation Study of Encoder
We conduct an ablation study where we train two variants of the encoder in Cond-FiP described as follows:
  • Cond-FiP (LIN): We sample SCMs with linear causal mechanisms during training of the encoder.
  • Cond-FiP (RFF): We sample SCMs with non-linear causal mechanisms during training of the encoder.
Note that for the training the subsequent decoder, we sample SCMs with both linear and rff causal mechanisms as in the main results ( Table 1Table 2, and Table 3). Note that in the main results, the encoder was trained by sampling SCMs with both linear and rff functional relationships. Hence, this ablation helps us to understand whether the strategy of training encoder on mixed functional relationships can bring more generalization to the amortization process, or if we should have trained encoders specialized for linear and non-linear functional relationships.
We present our results of the ablation study for the task of noise prediction, sample generation, and interventional generation in Table 19Table 20Table 21 respectively. Our findings indicate that Cond-FiP is robust to the choice of encoder training strategy! Even though the encoder for Cond-FiP (RFF) was only trained on data from non-linear SCMs, its generalization performance is similar to Cond-FiP where the encoder was trained on data from both linear and non-linear SCMs.